# Segment Any Mesh: Zero-shot Mesh Part Segmentation via Lifting Segment Anything 2 to 3D

### v1

@@ Results
Figure 4: A comparison of mesh segmentations produced by varying modalities, from left to right, untextured rendering, rendered surface normals only, rendered surface normals combined with rendered ShapeDiam scalars, and ShapeDiam scalars only.
CoSeg Dataset
Table 3 shows we exceed the performance of ShapeDiam on the CoSeg dataset in almost all metrics. Figure 5 shows one SAMesh output for each class of CoSeg.
Metric ShapeDiam SAMesh
Cut Discrepancy 0.391 0.361
Hamming Distance 0.164 0.164
Hamming Distance (Rf) 0.1 0.087
Hamming Distance (Rm) 0.228 0.241
Rand Index 0.217 0.21
Local CE 0.054 0.051
Global CE 0.093 0.08
Table 3: Princeton Mesh Segmentation Benchmark metrics for ShapeDiam vs SAMesh on CoSeg.
Figure 5: Examples of SAMesh outputs on the CoSeg dataset.
Princeton Mesh Segmentation Benchmark
The Princeton Mesh Segmentation Benchmark uses the mode number of human segmentation regions as the target number of GMM clusters, $k$, for ShapeDiam. For fairness, only for the Princeton Mesh Segmentation Benchmark, we search $\lambda$—the weight of the cost for alpha graph expansion. We select the smallest $\lambda$ that results in the number of segments produced by our method being within a predefined error margin of the mode value. This approach recognizes that while our method automatically determines the number of segments, it tends to focus on surface segmentation, whereas human annotations on the Princeton Mesh Segmentation Benchmark are often more component-oriented, so we usually increase lambda (more smoothing so fewer segmentations) compared to the default value.
Metric ShapeDiam SAMesh
Cut Discrepancy 0.278 0.3
Hamming Distance 0.163 0.175
Hamming Distance (Rf) 0.142 0.143
Hamming Distance (Rm) 0.184 0.207
Rand Index 0.173 0.186
Local CE 0.081 0.081
Global CE 0.126 0.128
Table 4: Princeton Mesh Segmentation Benchmark metrics for ShapeDiam vs SAMesh.
Metric ShapeDiam SAMesh
Cut Discrepancy 0.299 0.268
Hamming Distance 0.174 0.157
Hamming Distance (Rf) 0.157 0.137
Hamming Distance (Rm) 0.189 0.176
Rand Index 0.179 0.163
Local CE 0.088 0.076
Global CE 0.136 0.118
Table 5: Princeton Mesh Segmentation Benchmark metrics for ShapeDiam vs SAMesh without pathological categories
@@
Figure 6: Examples of SAMesh outputs on the Princeton Mesh Segmentation Benchmark dataset.
Table 4 shows that we perform comparably to ShapeDiam, albeit with better consistency metrics, since our method tends to produce cleaner boundaries. We identify 4 pathological classes (table, airplane, bird, mech). These are classes where our method’s surface-delineated segmentations differ significantly from more component-based segmentations. We argue the preferred result is often subjective and context dependent, but our segmentations can be more easily combined into component segmentations with manual intervention than vice versa. Examples of SAMesh outputs on these classes are shown in Figure 7. We show in Table 5, without considering these classes, our method exceeds that of ShapeDiam in performance.
Figure 7: SAMesh outputs on pathological classes (table, bird, mech, airplane) of the Princeton Mesh Segmentation Benchmark. We see that surfaces are prioritized over components, even with a $\lambda$ search. The ground truth and ShapeDiam segmentations for these classes can be view on (Chen, Golovinskiy, and Funkhouser 2009)’s website
Limitations
There are scenarios where parameter adjustments are necessary. Moreover, SAMesh may underperform when the underlying SAM model does not yield accurate results. Another current limitation is the evaluation time of SAM2, which serves as a bottleneck. The SAMesh process takes between 30 seconds to 1.5 minutes per mesh on a 30-core machine, much slower than that of ShapeDiam and inference of learning-based models. This could be improved by parallelizing SAM operations across multiple GPUs.
Conclusion
In this paper, we present SAMesh, a novel zero-shot approach to mesh part segmentation, leveraging multimodal rendering and 2D-to-3D lifting. By design, SAMesh seamlessly benefits from any improvements to 2D segmentation. This benefit doubles as a bottleneck to SAMesh’s effectiveness which is dependent on the underlying SAM model’s performance. We also note the relative lack of objective 3d segmentation metrics. While such a task is fundamentally subjective, we argue the more useful interface leans on the side of more segmented parts rather than fewer.
For future work, an interface could be developed to allow human users to refine the masks using 2D views, enabling them to perform tasks such as prompting, smoothing faces, and other SAM-based modifications. This approach could enhance practical applications like object understanding, texturing, and topology in cases where the method works well. For more precise tasks, human intervention could be integrated into the process to ensure accuracy. Additionally, we can leverage this human-in-the-loop process to curate a dataset and distill the dataset into a MeshCNN model (Hanocka et al. 2019), analogous to SAM’s data engine (Kirillov et al. 2023).

### v2

@@ 4 Experiments
Figure 6: Comparison of mesh part segmentations produced by varying input modalities to Segment Anything. From left to right, rendered surface normals only, rendered ShapeDiam scalars only, untextured rendering, and rendered surface normals combined with rendered ShapeDiam scalars (best).
Traditional Benchmarks
We now show Segment Any Mesh performs as well as ShapeDiam on traditional benchmarks, which ShapeDiam is overfitted to. Note that metrics should only be taken as an estimate of segmentation quality since mesh part segmentation can be quite subjective. Table 3 shows we match the performance of ShapeDiam, and Figure 5 shows example outputs of our method on the CoSeg dataset. Table 4 shows we also exceed the performance of ShapeDiam, and Figure 7 shows example outputs of our method on the Princeton Mesh Segmentation benchmark.
@@
Figure 7: Qualitative examples of Segment Any Mesh vs ShapeDiam outputs on the Princeton Mesh Segmentation Benchmark dataset. We provide additional comparisons with ShapeDiam in the supplemental.
Metric ShapeDiam Ours Cut Discrepancy 0.39 0.37 Hamming Distance 0.16 0.18 Hamming Distance (Rf) 0.10 0.08 Hamming Distance (Rm) 0.23 0.28 Rand Index 0.21 0.22 Local Consistency Error 0.06 0.05 Global Consistency Error 0.09 0.08
Table 3: Quantitative results for ShapeDiam vs Segment Any Mesh on the CoSeg dataset for automatic determination of the number of segmentation regions.
Metric ShapeDiam Ours Cut Discrepancy 0.34 0.31 Hamming Distance 0.21 0.17 Hamming Distance (Rf) 0.17 0.17 Hamming Distance (Rm) 0.26 0.17 Rand Index 0.22 0.21 Local Consistency Error 0.10 0.07 Global Consistency Error 0.16 0.12
Table 4: Quantitative results for Princeton Mesh Segmentation Benchmark for ShapeDiam vs Segment Any Mesh for automatic determination of the number of segmentation regions.
Additionally, the Princeton Mesh Segmentation Benchmark provides mesh segmentations for ShapeDiam that use the mode number of human segmentation regions as the target number of GMM clusters, $k$. We benchmark our method against these mesh segmentations to reflect the scenario when the target number of segmentation regions is roughly known. Specifically, we search $\lambda$—the weight of the cost for alpha graph expansion. We select the smallest $\lambda$ that results in the number of segments produced by our method within a predefined error margin of the mode value. We set $\tau_{R}=0.35$ while $\lambda$ is variable in the range $[1,15]$. Table 5 shows our method matches the performance of ShapeDiam.
Metric ShapeDiam Ours Cut Discrepancy 0.28 0.30 Hamming Distance 0.16 0.17 Hamming Distance (Rf) 0.14 0.14 Hamming Distance (Rm) 0.18 0.20 Rand Index 0.17 0.18 Local Consistency Error 0.08 0.08 Global Consistency Error 0.13 0.13
Table 5: Quantitative results for Princeton Mesh Segmentation Benchmark for ShapeDiam vs Segment Any Mesh given a reference number of segmentation regions determined by labelers.
5 Conclusion
Segment Any Mesh performs well on existing benchmarks and generalizes for diverse classes of meshes. This is due to our method being based on lifting rather than learning from limited diversity segmentation data or local shape descriptors. Lifting tackles the problem from a 2D view perspective, aligning with humans when analyzing object parts and affordances.
In the future, an interface can be developed refining the segmentations, which can further improve quality for targeted tasks such as object understanding, texturing, and retopology. We can also leverage this human-in-the-loop process to curate and distill a labeled segmentation dataset into a MeshCNN model [9].
Supplementary Material
Appendix A Additional Details for Human Evaluation Study
We use AWS Sagemaker as an evaluation interface. For the Segment Any Mesh versus shape diameter function comparison, we render a video of the meshes side by side, randomized, and ask the participants to rank which segmentations they prefer in order of (a), (b), (c), (d) where the letter corresponds to the respective segmentation index in the video. Below are the interfaces participants interacted with
Figure S1: Interace for Segment Any Mesh vs Shape Diameter Function experiment.
Figure S2: Interface for Segment Any Mesh vs Shape Diameter Function experiment. AWS Sagemaker does not support text input, so we resort to 24 labels corresponding to the ranking permutations.
We also include examples of frames from videos shown to the participants
Figure S3: Frames from a video shown to participants for Segment Any Mesh vs Shape Diameter Function preference experiment.
Figure S4: Frames from a video shown to participants for SAM input modality ablation experiment.
Appendix B Additional Visualizations
We also provide Segment Any Mesh vs Shape Diameter Function comparisons for CoSeg and Princeton Mesh Segmentation Benchmark.
Figure S5: Segment Any Mesh vs Shape Diameter Function on the CoSeg dataset.
Figure S6: Additional visualizations for Segment Any Mesh vs Shape Diameter Function on the CoSeg dataset.
Figure S7: Segment Any Mesh vs Shape Diameter Function on the Princeton Mesh Segmentation Benchmark for automatic determination of the number of segmentation regions i.e. not given the reference number of segmentation regions $k$.
Figure S8: Additional visualizations for Segment Any Mesh vs Shape Diameter Function on the Princeton Mesh Segmentation Benchmark for automatic determination of the number of segmentation regions.